Determine the area under the standard normal curve that lies between (a) Z=-0.79 and Z-0.79, (b) Z=-1.03 and Z-0, and (c) Ζ--0.63 and Z-032. (a) The area that lies between Z-079 and Z-079 is LJ (Round to four decimal places as needed.) (b) The area that lies between Z1.03 and Z-0 is (Round to four decimal places as needed.) (e) The area that lies between Z -063 and Zs 0.32 is Round to four decimal places as needed.)
Answer #1

(a)  

P[-0.79<Z<0.79]

ΡΙΖ < 0.79-PİZ <-0.79

=0.7852-0.2147.................................by using normal probability table.

=0.5705

b)

P[-1.03<Z<0]

=P[Zleq 0]-P[Zleq -1.03]

=0.5-0.1515..........................by using normal probability table.

=0.3485

c)

P[-0.63<Z<0.32]

PIZ < 0.321-PlZ <-0.63

=0.6255-0.2643..........................by using normal probability table.

=0.3612